The patient/client is to receive Nembutal sodium (pentobarbital sodium) 120 mg IM at bedtime prn. The medication is available as Nembutal sodium (pentobarbital sodium) 50 mg/mL. How many milliliters of medication would the patient/client receive?
Answer #1

ANSWER.

* Calculation of amount of Nembutal sodium required.

          Formula = Desired dose / Dose in hand * Quantity.

            Here

             Desired dose = 120 mg.

               Dose in hand = 50 mg.

                    Quantity = 1 ml.

      As per formula

        Amount of Nembutal sodium needed = D/H * Q

                                                                         = 120 / 50 * 1

                                                                         = 2.4 * 1

                                                                         = 2.4 ml.